I have a stock of 3.15mg/ml and I need to make a solution with water total volume of 200 ul with concentration of 200 ng/ml. I get serial dilution math but how do I make this pipettable volume?


1
Expert's answer
2022-01-20T00:28:01-0500

Given, M1 =3.15"\frac{mg}{mL}"  , V1 = ? 

M2 =200"\frac{mg}{mL}"  , V2 = 200"\mu L"

M1 in ng per mL

We know, 1 mg = 1.0 x 106 ng

So, 3.15 mg = 3.15 x 106 ng

Dilution law,   

M1V1=M2V2

V1=0.0127"\mu L"


We need to take 0.0127  of 3.15"\frac{mg}{mL}"   and diluted to 200"\mu L"  to prepare  of 200"\frac{mg}{mL}"of 200"\mu L"